A 1.0-kg ball with a speed of 3.0 m/s strikes a 2.0-kg stationary ball. The collision is completely inelastic. (a) Calculate the speed of the balls after collision. (b) Calculate the total kinetic energy change during the collision
Answer #1


the speed of the balls after collision v = m1*u1/(m1+m2) = 1*3/(2+1) = 1 m/sec...

change in K.E = Kf-Ki = (1/2)(m1+m2)*v^2 - (1/2)*m1*u1^2 =
(1/2)(1+2)*1^2-(1/2)*1*3^2 = 1.5-4.5 = -3 J.....